A 2.0 mm thick sheet of shielding allows 80% of incident gamma radiation to pass through it.
How many layers of this shielding are required such that less than 30% of incident gamma rays to pass through all layers? (to 2 s.f)
1
Expert's answer
2020-05-28T11:48:31-0400

For "\gamma" rays


"I=I_0e^{-\mu x}"


In our case


"0.8I_0=I_0e^{-\mu\cdot 0.002} \to \mu=111.57 m^{-1}"


So,


"0.3I_0=I_0e^{-111.57\cdot x} \to x=111.57 m^{-1}\approx0.011m"


"n=\frac{0.011}{0.002}=5.5\approx6"


6 plates are required.